yomomalikebracobama:

Solve 5c − c + 10 = 34. (1 point) a −6 b 6 c 11 d −11

AZ:

Do you know how to add/subtract like terms? If you have 5 apples and you give away one apple, then how many apples are you left with?

yomomalikebracobama:

3

yomomalikebracobama:

4****

AZ:

Yes so basically 5c - c is like 5 apples minus 1 apple you would be left with 4 apples or 4c

AZ:

So to simplify your question, we now have 4c + 10 = 34 We want to move that 10 to the other side so that way we can get 'c' all by itself. Since we're adding 10 on the left side, we need to do the opposite of addition. What is the opposite of addition?

yomomalikebracobama:

-10 on both sides

yomomalikebracobama:

right?

AZ:

Exactly! And that way the 10 cancels out on the left side of the equal sign 4c + 10 - 10 = 34 - 10 can you simplify it now? 10 - 10 is 0 what is 34 - 10 = ?

yomomalikebracobama:

24

yomomalikebracobama:

or -24 i canr remeber i havent been focused all day

AZ:

positive 24 if you had 34 dollars and you gave me 10 dollars, how much would you be left with?

yomomalikebracobama:

24 cuz i wouldnt give u any money

yomomalikebracobama:

34*

AZ:

you got jokes, nice

yomomalikebracobama:

lol but i would have 24

AZ:

anyway, so now we have 4c = 24 what number multiplied by 4 gives you 24? alternatively, you can divide 4 on both sides to find out what c is equal to

yomomalikebracobama:

idk lemme as google

yomomalikebracobama:

idk lemme as google

AZ:

so 4c is the same thing as 4 multiplied by c 4 * c = 24 if we want to move 4 to the other side of the equal sign, we have to do the opposite of multiplication on both sides. That means, we have to divide 4 on both sides

yomomalikebracobama:

6

AZ:

You got it!

yomomalikebracobama:

c=6

AZ:

That's your answer!

yomomalikebracobama:

i got 80 out of 100

yomomalikebracobama:

cuz SOMBOSY IN HERE GAVE ME WRONG AWNSER

yomomalikebracobama:

i shall hunt him down

AZ:

If you want to post your other questions in a separate post, we can try to figure out which one you got wrong

yomomalikebracobama:

ok bet